[Brahms-l] Fw: SPY news update: Major change of plan ... Time meeting on Tuesday (Apr. 20)

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Mon Apr 19 2004 - 16:48:43 EDT
Well,
It is hard to see how to plan for this run; try to digest the following.
I will circulate another shift schedule shortly - but stick to the owl-shift
until Friday

Flemming


----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106
----- Original Message ----- 
From: "Kin Yip" <kinyip@bnl.gov>
To: <kinyip@bnl.gov>
Sent: Monday, April 19, 2004 4:37 PM
Subject: SPY news update: Major change of plan ... Time meeting on Tuesday
(Apr. 20)


> ====== Scheduling Physicist -YIP (ie. SPY) news channel ========
>
> It's a confusing day and I've told people different plans a couple times
> within
> today.   I can only apologize :-)  But we still have to CHANGE !
>
> **************************************************************************
>
> (1) Schedule (as I have told major parties):
>
>  (A) There will NO AGS access on Tuesday morning !
>   ** The AGS polarimeter work is tentatively postponed to next Tuesday
>       (Apr. 27). They have to change the silicons at some point and
>       it's difficult to postpone it again.
>
>  (B) There will be an OWL shift this coming midnight (~00:01-08:00, Apr.
> 20).
>
>  (C) It looks like that the one week data taking is now postponed to start
>        on Friday, Apr. 23 (though there is daily re-evaluation).
>
>  (D) In principle, we'll have an OWL shift every night until Friday but
the
>         machine people may take over those OWL shifts if it's necessary.
>
> **************************************************************************
>
> (2) There will be a Time meeting this Tuesday (Apr. 20) at 13:30 in the
>       Synder Room (911A).
>
> **************************************************************************
>
> (3) Status :
>
> For the polarized-proton program in RHIC,  the main
> concern at this moment is that we don't have reliable polarization
> in the yellow beam (often/sometimes consistent with 0, within error).
>
> Another goal is to get higher and higher luminosity at physics stores.
>
> The polarization and beam intensity in the blue ring are probably OK
> with the Jet target alone (which needs only the blue beam) but it's
> probably
> not good enough (eg. 0% polarization in the yellow ring) for the other
> experiments to make use of.
>
> Today is not a good day for RHIC (one machine problem after another) but
> it seems
> that NSRL is doing the last expt. of the day, right on schedule, even
> though there
> was a 1.5 hour down-time.    NSRL has the last 600 MeV Fe run today and
> tomorrow NSRL will run 1 GeV Ti.  This run of NSRL will finish this
> Thursday
> (Apr. 22).
>
> ====== ======= ====== ======= ====== ======= ====== =======
>
> This SPY news channel solicits YOUR input of any sorts.
> You'll be paid back with loads of gratefulness :-)
>
> Let me know if you want to get out of this private email mailing list.
>
> SPY's webpage :
>
> http://www.c-ad.bnl.gov/kinyip/SchedPhys/
>
> =================================================================
>
> Disclaimer: Whatever is being said here is not always 100% accurate
> though I try my best to avoid being 100% inaccurate :-)
>
>


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Apr 19 16:41:58 2004

This archive was generated by hypermail 2.1.8 : Mon Apr 19 2004 - 16:42:26 EDT